OpenStudy (anonymous):

prove that arctan x+arctan y= arctan x+y/1-xy

OpenStudy (anonymous):

Can you put parentheses where they belong toward the end of the equation?

OpenStudy (anonymous):

is 1-xy over arctan x+y or just under y?

OpenStudy (anonymous):

hitthis part of the question i have answered i am struggling with the second part which is slightly tricky. it says use this relationship to show that if arctan (x) + arctan (y) + arctan(z) = pi/2 then xy + yz + zx = 1. i

OpenStudy (anonymous):

the relationship is arctan (x) + arctan (y)= arctan (x+y)/(1-xy)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

I'm sorry, but I don't think I can help you with that one. Is it supposed to be a system of equations?

OpenStudy (anonymous):

i am struggling with simplifying the terms

OpenStudy (anonymous):

For the first or second one?

OpenStudy (anonymous):

i sthe second one as it will follow the same form as the first one but you have the extra term added arctan z. this is how far i have got

OpenStudy (anonymous):

(x+y+z-xyz)/(1-xy-xz+yz)=pi/2